### v1

@@ 1 Introduction
Ultra-high-energy (UHE) photons of astrophysical origin, with energies above $10^{18}$ eV, are of interest for two main reasons. First, they can be produced via the Greisen-Zatsepin-Kuzmin (GZK) mechanism [greisen1966end, Zatsepin:1966jv, Wdowczyk:1971jk, Michalak:1990iw, Gelmini:2007jy, Hooper:2010ze, Gelmini:2022evy] in the interstellar medium. The resulting photon flux depends on the mass composition of cosmic rays, thus providing an independent tool for their analyses. Second, registering ultra-high-energy photons beyond the expected GZK flux level would provide strong evidence for models beyond the Standard Model of particle physics [berezinsky1998signatures, Anchordoqui:2021crl, Fairbairn:2009zi], including, in particular, decaying dark matter [Berezinsky:1997hy, Kuzmin:1997jua, Kuzmin:1999zk, Kalashev:2016cre, Kalashev:2020hqc, Chianese:2021jke, Das:2023wtk] and violation of Lorentz symmetry [coleman1999high, galaverni2008lorentz, rubtsov2014prospective, PierreAuger:2021tog].
@@
Experimental searches for ultra-high-energy photons rely on complementary detection techniques, each with distinct advantages and limitations. Fluorescence Detectors measure fluorescence light from extensive air showers (EAS), allowing for the direct observation of the longitudinal shower development and the depth of shower maximum ($X_{max}$). This provides excellent discrimination between photon and hadron primaries, but their operation time is restricted to clear, moonless nights, resulting in a limited duty cycle that constrains the exposure accumulation for ultra-high-energy events. Conversely, scintillator-based Surface Detectors, measuring secondary particles at the ground level, operate continuously, offering the vast exposure required to probe low-flux regimes. However, they face challenges in distinguishing muon and electromagnetic components of EAS, whose ratio is the key for distinguishing hadron-induced and photon-induced EAS. To address this, experiments like Yakutsk [Glushkov:2009tn] employed dedicated muon detectors to suppress the hadronic background. Similarly, the Pierre Auger Observatory has set stringent limits in the Southern Hemisphere by utilizing both hybrid observations [PierreAuger:2024ayl] and water-Cherenkov Surface Detectors [abreu2023search], which possess different sensitivities to the muon component compared to scintillators. Early constraints were also established by the AGASA experiment [Shinozaki:2002ve]. None of the experiments reported an excess of photon-induced events above the expected hadronic background, producing photon-like events due to fluctuations of EAS evolution.
In this paper, we search for ultra-high-energy photons using Telescope Array (TA) experiment data. TA is a large-scale cosmic ray experiment located in Utah, USA, 39°17’ 49” N 112°54’ 31” W, at approximately 1.4 km above sea level [abu2012surface, tokuno2012new, udo2017telescope]. It aims to study ultra-high-energy cosmic rays (UHECR) using its Surface Detector (SD), Fluorescence Detector (FD), and Low Energy (TALE) extension. Operating since 2008, TA has accumulated a large number of EAS events, allowing for thorough studies of cosmic ray properties.
@@
Unlike FD, SD operates under almost all weather conditions, but provides less informative data. To compensate for this, we employ a neural network classifier that is capable of leveraging complex patterns in data to effectively discern between photon- and proton-induced events. In the astrophysical field, neural networks have been successfully applied in TA [Kalashev:2020fzg, Ivanov:2020nfo, Kalashev:2021vop], Pierre Auger Observatory [PierreAuger:2021fkf, PierreAuger:2023gmj, Glombitza:2020yhw], KASCADE-Grande [Kuznetsov:2023pvo], and neutrino telescopes [Huennefeld:2017pdh, IceCube:2021umt, DeSio:2019lcr, Spisso:2022tpm, Kharuk:2022abi, Kharuk:2023xnl]. These examples illustrate the potential of using machine learning in our analysis using TA SD data.
@@
Our method incorporates a blind optimization of the neural network classification threshold to establish the most stringent constraint on the photon flux. To mitigate discrepancies between Monte Carlo simulations and the experimental data, the technique includes a procedure for fine-tuning the neural networks using a relevant subset of the experimental data. This allows us to achieve robust and efficient separation between hadron-induced and photon-induced EAS.
@@
The TA SD array, depicted in Figure 2, comprises 507 above-ground scintillation stations arranged in a square grid covering an area of approximately 700 km2. Each SD station consists of two layers of plastic scintillator, which register signals with a time resolution of 20 ns [abu2012surface]. All stations are calibrated in real time using atmospheric muons, with the signal measured in conventional units called a "minimal ionizing particles" (MIPs). For our analysis, we retain 128 consecutive time bins from each layer of triggered SD stations, which we refer to as waveforms. An example of a waveform registered by a single SD station is presented in Figure 2.
Figure 1: Schematic representation of the arrangement of the TA SD and FD stations, [abu2012surface].
@@ 2 Method
1. Data selection and preparation. We perform Monte Carlo (MC) simulations of 14 years of TA SD data, modeling air showers initiated by protons and photons. Heavy nuclei such as iron produce showers that develop earlier, exhibit denser cores, and higher muon-to-electromagnetic ratios than proton-induced cascades [Ros:2011zg, ros2013improving, cazon2023universality]. This makes them easier to distinguish from photon-induced air showers, yielding protons as the main background source to photons signal.
@@
3. Blind optimization of the classification threshold. At present, the most energetic photons observed in the experiment have energies of the order of PeV [TibetASgamma:2021tpz, LHAASO:2021gok]. Strong constraints have been set on the upper limit of photon flux at EeV energies [Fomin:2017ypo, KASCADEGrande:2017vwf, PierreAuger:2025jwt, aab2017search, abreu2023search, TelescopeArray:2018rbt] without detecting photon signal — the number of observed photon candidates was consistent with background expectations. Based on this, we optimize the neural network’s classification threshold to provide the strongest constraint on the photon flux after unblinding the data and calculating the actual number of photon candidates.
@@
We minimize this value by performing blind optimization of the classification threshold on the MC data set assuming HiRes [HiRes:2007lra] and $E^{-2}$ differential energy spectra for proton and photon primaries, respectively.
@@
The estimation of the numerator in Equation 2.1 requires a multi-step procedure. Initially, the proton suppression level, $P(\xi)$, is quantified on the MC sample. It is defined as the ratio of events incorrectly identified as photons to the total number of proton events for a given $\xi$. Subsequently, the expected number of false photons due to the proton background in the experimental data, $n^{\text{bg}}(\xi)$, is estimated by multiplying $P(\xi)$ by the total number of experimental events. Further, the corresponding $\mathbf{95\%}$ confidence level upper limit is determined using the Feldman-Cousins statistics [Feldman:1997qc] under the assumption of zero background, which is designated as $\sigma^{95}(n^{\text{bg}}(\xi))$. Given that event registration is inherently a random process, $\mathbf{n^{\text{bg}}(\xi)}$ should be considered as a Poisson random variable, hence optimizing $\sigma^{95}(n^{\text{bg}}(\xi))$ expectation value.
@@ 3 Data sets and Monte Carlo simulations
MC simulations conducted for this study encompass the complete evolution of air showers and the subsequent response of TA SD stations. These simulations are performed to model 14 years of TA operation, using CORSIKA [Heck:1998vt] for air shower simulations and GEANT4 [GEANT4:2002zbu] for detector response modeling. The simulations adhered to the actual real-time calibration tables of the TA SD to ensure an accurate representation of the detector behavior.
@@
For proton-induced events, we employed three high-energy hadronic interaction models: QGSJET-II-04 [ostapchenko2006qgsjet], EPOS-LHC [pierog2009epos], and Sibyll 2.3d [ahn2009cosmic]. Low-energy hadronic and electromagnetic interactions were simulated using FLUKA [ferrari2005fluka] and EGS4 [nelson1985egs4], correspondingly.
@@
The simulations were performed in three energy ranges, starting from $10^{17.45},~10^{18.95},$ and $10^{19.45}$ eV. Within each sample, the proton energy spectrum is modeled to follow the one measured by the HiRes experiment [HiRes:2007lra]. Due to the sharp drop-off of the spectrum at high energies, such splitting is required for a comprehensive comparison of the MC and experimental data at different energy scales. These datasets were subsequently used for setting upper limits on photon flux at different energy cutoffs. The number of simulated air showers is the same for all of the high-energy hadronic interaction models.
@@
Photon-induced events are simulated using the QGSJET-II-04, FLUKA, and EGS4. We use the PRESHOWER code [Homola:2003ru] to account for geomagnetic interactions resulting in magnetic $^{+}ee^{-}$ pair production. The corresponding air showers are dominated by their electromagnetic component, hence the effect of choosing a particular high-energy hadron interaction model should be negligible. We test this assumption with Monte Carlo simulations in Section 5. The differential energy spectrum is chosen to be $E^{-1}$ to cover the whole range of energies within a single data set, thus avoiding underrepresentation of high-energy events.
@@
MC and experimental data were passed through a reconstruction procedure [TelescopeArray:2013yze] optimized to infer the parameters of photon-induced events. We estimate the energy of events based on the look-up table constructed using MC photon data set [TelescopeArray:2018rbt], which is a function of the reconstructed $S_{800}$ value (scintillator signal density at the distance of 800 m from the core), zenith and azimuth angles. Thus obtained energy corresponds to average energy of the primary photon, inducing a shower with the same arrival direction and $S_{800}$. The full list of reconstructed features, presented in Appendix A, includes geometrical and primary particle mass-sensitive parameters, such as reconstructed zenith angle and Linsley front curvature.
@@
The chosen set of event quality cuts is closer to the standard ones used for anisotropy studies with proton primaries in TA. It provides a balance between reconstruction accuracy and photon exposure. Note that the cut in the zenith angle is slightly stricter than in our previous studies [TelescopeArray:2018rbt, TelescopeArray:2019mwd]. Additionally, since it is impossible to properly simulate the signal in the saturated stations, we exclude them both from MC-simulated and experimental data.
To create a balanced data set for neural network training and evaluation, we employ a mixing procedure for proton-induced and photon-induced events. The energy range is divided into bins of constant width ($0.05$ in $\text{log}_{10}$ scale), in each of which the events are randomly sampled so that the total number of registered and successfully reconstructed events of both types is equal. This makes the energy spectrum of proton- and photon-induced events the same, which mitigates potential biases related to event energy, which could otherwise be exploited by the neural network as a discriminative yet unphysical parameter. The resulting energy spectrum of events passing event quality cuts in MC sample is presented in Figure 3.
@@
Figure 3: Energy distribution of the MC sample used for neural network training with mixed photon-induced and proton-induced events after applying event selection criteria.
For the experimental data, we have excluded all lightning-correlated events. That is, we reject events registered within a 10-minute interval related to lightnings recorded by the National Lightning Detection Network [cummins2009overview] at the location of the TA SD. This step is necessary as lightning can be registered by TA SD as photon-like events [abbasi2017bursts].
@@ 4.1 Architecture
The neural network architecture used in this study is a slight modification of the one used previously for the analysis of mass composition [kalashev2022deep] and the estimation of the flux of photons [Kharuk:2023gts]. Its architecture, presented in Figure 4, consists of three main components, each focusing on different aspects of the cosmic ray-induced air shower:
@@
The splitting of station activation time into two components, items 4 and 5 in the list above, is useful as it provides the neural network the information on the front curvature. The data for the stations that were not triggered in an event are filled with zeros.
@@
    First, waveforms of all SD station are passed through a waveform encoder, which extracts features representing waveform characteristics using a combination of convolutional and recurrent layers. In particular, we use LSTM initiator sub-network to initialize the state of the recurrent layers, utilizing reconstructed event properties, station coordinates and activation times (see Fig. 4).
@@
Figure 5: Architecture of the temporal station bundle. See the temporal station bundle in Figure 4 for layers characteristics.
@@
In total, the neural network has 29000 trainable parameters. We used the Adam optimizer [kingma2014adam] and employ early stopping for terminating neural network training. The neural network is implemented using the TensorFlow library [tensorflow2015-whitepaper].
@@ 4.2 Training Process
(2) We used a special loss function, namely, focal loss [lin2017focal]:
@@
Monte Carlo simulations inevitably involve simplifications, such as the application of thinning and de-thinning procedures [Stokes:2011wbu], as well as uncertainties related to the modeling of air shower evolution and the response of detector stations. To account for such imperfections, and avoid overfitting to MC data as opposed to experimental data, we employed two complementary strategies.
@@
Using a burn sample to fine-tune the neural network constitutes the novelty of our approach. It helps the network adapt to the nuances of real data that may not be fully captured in simulations. Importantly, we observed that the effect of using burn sample on neural network metrics is negligible, indicating that there are no major MC-experimental data discrepancies. This step was inspired by surrogate supervision and labeling [Tajbakhsh:8759553], where a limited subset of data with approximate or indirectly obtained annotations is employed to guide model adaptation to new domains.

### v2

@@ 1 Introduction
Cosmogenic ultra-high-energy (UHE) photons, with energies above $10^{18}$ eV, are of interest for two main reasons. First, they can be produced via the Greisen-Zatsepin-Kuzmin (GZK) mechanism [44, 85, 84, 69, 40, 47, 41] in the interstellar medium. The resulting photon flux depends on the mass composition of cosmic rays, thus providing an independent tool for their analyses. Second, registering ultra-high-energy photons beyond the expected GZK flux level would provide strong evidence for models beyond the Standard Model of particle physics [26, 23, 35], including, in particular, decaying dark matter [25, 63, 64, 53, 55, 30, 33] and violation of Lorentz symmetry [31, 39, 76, 16].
@@
Experimental searches for ultra-high-energy photons rely on complementary detection techniques, each with distinct advantages and limitations. Fluorescence detectors measure fluorescence light from extensive air showers (EAS), allowing for the direct observation of the longitudinal shower development and the depth of shower maximum ($X_{max}$). This provides excellent discrimination between photon and hadron primaries, but their operation time is restricted to clear, moonless nights, resulting in a limited duty cycle that constrains the exposure accumulation for ultra-high-energy events. Conversely, scintillator-based surface detectors, measuring secondary particles at the ground level, operate continuously, offering the vast exposure required to probe low-flux regimes. However, they face challenges in distinguishing muon and electromagnetic components of EAS, whose ratio is the key for distinguishing hadron-induced and photon-induced EAS. To address this, experiments like Yakutsk [43] employed dedicated muon detectors to suppress the hadronic background. Similarly, the Pierre Auger Observatory has set stringent limits in the Southern Hemisphere by utilizing both hybrid observations [11] and water-Cherenkov Surface Detectors [15], which possess different sensitivities to the muon component compared to scintillators. Early constraints were also established by the AGASA experiment [77].
A different approach was employed in the Telescope Array (TA) experiment. TA is a large-scale cosmic-ray experiment located in Utah, USA, 39°17’ 49” N 112°54’ 31” W, at approximately 1 400 m above sea level [18, 82, 83]. It aims to study ultra-high-energy cosmic rays (UHECR) using its Surface Detector (SD) and Fluorescence Detector (FD). Since high exposure is crucial for probing low fluxes, TA used its SD data to set an upper limit on the photon flux using composition-sensitive parameters, such as EAS front curvature [17], which were further improved using boosted decision trees [6].
To date, no experiments have reported a significant excess of photon-induced events above the expected hadronic background (see Figure 9). This background arises from the imperfect discrimination between photon and hadron primaries, driven by the inherent stochasticity of air-shower evolution. Specifically, hadron-induced showers can fluctuate to develop a dominant electromagnetic component. A primary example is the production of a leading $\pi^{0}$ that decays via $\pi^{0}\rightarrow\gamma\gamma$ channel [72]; such events are virtually indistinguishable from primary photon showers and constitute an irreducible background.
@@
In this paper, we search for ultra-high-energy photons in 14 years of TA SD data using a neural network classifier, capable of leveraging complex patterns in the data to discern between photon-induced and proton-induced events. In the field of Astroparticle Physics, neural networks have been successfully applied in TA [56, 50, 54], the Pierre Auger Observatory [2, 12, 42], KASCADE [66, 65], and neutrino telescopes [49, 48, 34, 78, 57, 58], illustrating the potential of machine learning for our task.
@@
The present analysis is the natural evolution of [59], which demonstrated the viability of a neural-network-based photon search for TA SD data. The novelty of our approach is that we ensure that predictions of our neural network on the experimental data are reliable by mitigating the discrepancies between Monte Carlo simulations and the experimental data. For this purpose our technique includes a procedure for fine-tuning the neural networks using a relevant subset of the experimental data. Our method also incorporates a blind optimization of the neural network classification threshold to establish the most stringent constraint on the photon flux.
@@
The neural network operates directly on the signals recorded by the TA SD stations, whose main features we now summarize. The TA SD array, depicted in Figure 2, comprises 507 above-ground scintillation stations with 1.2 km spacing, arranged in a square grid covering an area of approximately 700 km2. Each SD station consists of two layers of plastic scintillator, which register signals with a time resolution of 20 ns [18]. All stations are calibrated in real time using atmospheric muons, with the signal measured in conventional units called a “minimal ionizing particles” (MIPs). For our analysis, we retain 128 consecutive time bins from each layer of triggered SD stations, which we refer to as waveforms. An example of a waveform registered by a single SD station is presented in Figure 2.
Figure 1: Schematic representation of the arrangement of the TA SD and FD stations, [18].
@@ 2 Method
1. Data selection and preparation. We perform Monte Carlo (MC) simulations of 14 years of TA SD data, modeling air showers initiated by protons and photons (see Section 3 for details). Heavy nuclei such as iron produce showers that develop earlier, exhibit denser cores, and higher muon-to-electromagnetic ratios than proton-induced cascades [75, 74, 29]. This makes them easier to distinguish from photon-induced air showers, yielding protons as the main background source to photons signal.
@@
3. Blind optimization of the classification threshold. At present, the most energetic photons observed in the experiment have energies of the order of PeV [22, 28]. Strong constraints have been set on the upper limit of photon flux at EeV energies [38, 24, 13, 1, 15, 6] without detecting photon signal — the number of observed photon candidates was consistent with background expectations. Based on this, we optimize the neural network’s classification threshold to provide the strongest constraint on the photon flux before unblinding the data and calculating the actual number of photon candidates.
@@
We minimize this value by performing blind optimization of the classification threshold on the MC data sets. For protons, we assume HiRes [4] energy spectrum, which is close to experimental estimates. For photon primaries we choose $E^{-2}$ differential energy spectrum, following common choice in similar studies [61, 11, 15].
@@
The estimation of the numerator in Equation 2.1 requires a multi-step procedure. Initially, the proton suppression level, $P(\xi)$, is quantified on the MC sample. It is defined as the ratio of events incorrectly identified as photons to the total number of proton events for a given $\xi$. Subsequently, the expected number of false photons due to the proton background in the experimental data, $n^{\text{bg}}(\xi)$, is estimated by multiplying $P(\xi)$ by the total number of experimental events. Further, the corresponding $\mathrm{95\%}$ confidence level upper limit is determined using the Feldman-Cousins statistics [36] under the assumption of zero background, which is designated as $\sigma^{95}(n^{\text{bg}}(\xi))$. Given that event registration is inherently a random process, $\mathbf{n^{\text{bg}}(\xi)}$ should be considered as a Poisson random variable, hence optimizing $\sigma^{95}(n^{\text{bg}}(\xi))$ expectation value.
@@ 3 Data sets and Monte Carlo simulations
MC simulations conducted for this study encompass the complete evolution of air showers and the subsequent response of TA SD stations. These simulations are performed to model 14 years of TA operation, using CORSIKA version 7.7420 [45] for air-shower simulations and GEANT4 [20] for detector response modeling. The simulations randomly sample event timestamps and reproduce the corresponding experimental conditions by using actual real-time calibration tables and list of non-functional TA SD stations, thus ensuring an accurate representation of the detector behavior.
@@
For proton-induced events, we employed three high-energy hadronic interaction models: QGSJET-II-04 [71], EPOS-LHC [73], and Sibyll 2.3d [21]. Low-energy hadronic and electromagnetic interactions were simulated using FLUKA [37] and EGS4 [70], correspondingly.
@@
The simulations were performed in three energy ranges, starting from $10^{17.45},~10^{18.95},$ and $10^{19.45}$ eV. Within each sample, the proton energy spectrum is modeled to follow the one measured by the HiRes experiment [4]. Due to the sharp drop-off of the spectrum at high energies, such splitting is required for a comprehensive comparison of the MC and experimental data at different energy scales. These datasets were subsequently used for setting upper limits on photon flux at different energy cutoffs. The number of simulated air showers is the same for all of the high-energy hadronic interaction models.
@@
Photon-induced events are simulated using the QGSJET-II-04, FLUKA, and EGS4. We use the PRESHOWER code [46] to account for geomagnetic interactions resulting in magnetic $^{+}ee^{-}$ pair production. The corresponding air showers are dominated by their electromagnetic component, hence the effect of choosing a particular high-energy hadron interaction model should be negligible. We test this assumption with Monte Carlo simulations in Section 5. The differential energy spectrum is chosen to be $E^{-1}$ to cover the whole range of energies within a single data set, thus avoiding underrepresentation of high-energy events.
@@
MC and experimental data were passed through a reconstruction procedure [17] optimized to infer the parameters of photon-induced events. Since we are searching for photon-like events, we estimate the energy of events based on the look-up table constructed using MC photon data set [6], which is a function of the reconstructed $S_{800}$ value (scintillator signal density at the distance of 800 m from the core), zenith and azimuth angles. Thus obtained energy, $E_{\gamma}$, corresponds to average energy of the primary photon, inducing a shower with the same arrival direction and $S_{800}$. In contrast to proton-optimized energy reconstruction, this approach ensures that we set corrects limits on the photon flux for a specified energy threshold. For proton primaries, $E_{\gamma}$ yields systematically higher energy estimate than $E_{p}$, an analogous unbiased energy estimator for the proton primaries. For instance, at $\sim 10$ EeV Monte-Carlo energy the mean $E_{\gamma}$ for proton primaries is $\sim 55\%$ higher than their $E_{p}$. The full list of reconstructed features, presented in Appendix A, includes geometrical and primary particle mass-sensitive parameters, such as reconstructed zenith angle and Linsley front curvature.
@@
The chosen set of event quality cuts is close to the standard ones used for anisotropy studies with proton primaries in TA, providing a balance between reconstruction accuracy and photon exposure. The cut on the reconstructed shower core position ensures that geometry and charge deposition profiles are well-understood, without extrapolating peak values outside TA SD array. Note that the cut in the zenith angle is slightly stricter than in our previous studies [6, 7]. Additionally, since it is impossible to properly simulate the signal in the saturated stations, we exclude them both from MC-simulated and experimental data.
To create a balanced data set for neural network training and evaluation, we employ a mixing procedure for proton-induced and photon-induced events. The energy range is divided into bins of constant width ($0.05$ in $\text{log}_{10}$ scale), in each of which the events are randomly sampled so that the total number of registered and successfully reconstructed events of both types is equal. The number of events in each bin is limited by data set with lower event count, up to sampling fluctuations. This makes the energy spectrum of proton- and photon-induced events the same, which mitigates potential biases related to event energy, which could otherwise be exploited by the neural network as a discriminative yet unphysical parameter. The resulting energy spectrum of events passing event quality cuts in MC sample is presented in Figure 3.
@@
Figure 3: Energy distribution of the MC samples after applying event selection criteria. Dotted lines show energy distributions of each individual data set. The combined sample, which as equal mixture of proton-induced and photon-induced events with the same energy spectrum, was used for neural network training.
For the experimental data, we have excluded all lightning-correlated events. That is, we reject events registered within a 10-minute interval related to lightnings recorded by the National Lightning Detection Network [32] at the location of the TA SD. This step is necessary as lightning can be registered by TA SD as photon-like events [9].
@@ 4.1 Architecture
The neural network architecture used in this study is a slight modification of the one used previously for the analysis of mass composition [52] and the estimation of the flux of photons [59]. Its architecture, presented in Figure 4, consists of three main components, each focusing on different aspects of the cosmic-ray-induced air shower:
@@
The splitting of station activation time into two components, items 3 and 4 in the list above, is useful as it provides the neural network the information on the front curvature. The data for the stations that were not triggered in an event are filled with zeros.
@@
    Waveforms from upper and lower layers of SD station form a two-channel one-dimensional “image”, which can also be considered as a sequence. We apply a waveform encoder to each of the registered waveforms to extract features representing waveform characteristics using a combination of convolutional and recurrent layers (specifically, long short-term memory (LSTM) blocks). We use LSTM initiator sub-network to initialize the state of the recurrent layers, utilizing reconstructed event properties, station coordinates and activation times (see Fig. 4).
@@
Figure 5: Architecture of the temporal station bundle. See the temporal station bundle in Figure 4 for layers characteristics. For simplicity, we show only one waveform per SD station.
@@
In total, the neural network has 29000 trainable parameters. We used the Adam optimizer [62] and employ early stopping for terminating neural network training. The neural network is implemented using the TensorFlow library [3].
@@ 4.2 Training Process
(2) We used a special loss function, namely, focal loss [68]:
@@
Monte Carlo simulations inevitably involve simplifications, such as the application of thinning and de-thinning procedures [79], as well as uncertainties related to the modeling of air-shower evolution and the response of detector stations. To account for such imperfections, and avoid overfitting to MC data as opposed to experimental data, we employed two complementary strategies.
@@
Using a burn sample to fine-tune the neural network constitutes the novelty of our approach. It helps the network adapt to the nuances of real data that may not be fully captured in simulations. Importantly, we observed that the effect of using burn sample on neural network metrics is negligible, indicating that there are no major MC-experimental data discrepancies. This step was inspired by surrogate supervision and labeling [80], where a limited subset of data with approximate or indirectly obtained annotations is employed to guide model adaptation to new domains.